OpenStudy (anonymous):

REPOSTED (Projectile Flight) A ball is rolled off the edge of a swimming pool and strikes the water 2.0 away 0.30 seconds later (a) what was the initial speed of the ball? (b) how high is the edge of the pool from the surface of water?

OpenStudy (anonymous):

You forgot to mention the unit for 2.0 meters? (A) Since the horizontal speed does not change during projectile motion use: \[V_x = \frac{x_f - x_i}{t_f - t_i}\] (B) You can use the distance equation for the vertical direction: \[Y_f = Y_i + V_it + \frac{1}{2}a_yt\] or \[D_y = V_it + \frac{1}{2}a_yt\] I would start with a picture with all motion problems and then put the y,v,t into the picture

OpenStudy (anonymous):

Do these equations make sense? Would you like to see a picture?

OpenStudy (anonymous):

I don't get the second equation. May I see a picture about it?

OpenStudy (anonymous):